OpenStudy (anonymous):

What is the value of work done on an object when a 70-newton force moves it 9.0 meters in the same direction as the force?

OpenStudy (souvik):

W=Fdcos (theta) theta is the angle between F(force) and d(distance)

OpenStudy (anonymous):

theta??

OpenStudy (souvik):

\[\theta \]

OpenStudy (souvik):

\(W=F*d*cos\theta\)

OpenStudy (anonymous):

thanks

OpenStudy (souvik):

np:)